News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 22 February 2002 10.00am EST

During the last week the observing schedule was completed as planned.

Chandra passed through eclipses 7, 8 and 9 of the 2002 Spring season on February 16, 19 and 21 with nominal power and thermal performance. A real-time procedure was carried out on February 21 to reconfigure the spacecraft's power subsystem for the short eclipse that will occur on February 24.

An image was released last week of the hot galaxy cluster 1E0657-56, revealing a bow-shaped shock wave believed to be the result a smaller group of galaxies merging with 1E0657-56. The image can be viewed at http://chandra.harvard.edu/photo/2002/0066/index.html.

The observing schedule for the next week is shown below. It includes a series of measurements of the Chandra Aspect Camera Assembly's dark current, and the third part of a target of opportunity, a black hole transient (XTE J1650-500), that was triggered in late January.

------------------------------------------
3C294                                   Feb 25
FTM0729+3336            ACIS-S NONE     Feb 26
ACA dark current
Radiation Belts
3C58                    HRC-S  NONE     Feb 27
3C294                   ACIS-S NONE
Q1416+067               ACIS-S NONE     Feb 28
Radiation Belts                         Mar 1
1202-262                ACIS-S NONE
XTE J1650-500           ACIS-S NONE
1E 1740.7-2942          HRC-I  NONE     Mar 2
M17                     ACIS-I NONE

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports